OronSH:

How do I solve this? a, b, c, d, e, f are positive numbers, if a+b+c+d+e+f=6 and a^2+b^2+c^2+d^2+e^2+f^2=36/5 how do I find the largest possible value of a^3+b^3+c^3+d^3+e^3+f^3?

cuzican:

@cuzican wrote:
( 4 a+b+c+d ​ ) 2 ≤ 4 a 2 +b 2 +c 2 +d 2 ​ (i)(Using Tchebycheff's Inequality) Where a+b+c+d+e=8 and a 2 +b 2 +c 2 +d 2 +e 2 =16 ∴ Eq. (i), reduces to ( 4 8−e ​ ) 2 ≤ 4 16−e 2 ​ ⇒ 64+e 2 −16e≤4(16−e 2 ) ⇒ 5e 2 −16e≤0 ⇒ e(5e−16)≤0 (Using number line rule) ⇒ 0≤e≤ 5 16 ​ Thus, range of eϵ[0, 5 16 ​ ] does this help? ( 4 a+b+c+d ​ ) 2 ≤ 4 a 2 +b 2 +c 2 +d 2 ​ (i)(Using Tchebycheff's Inequality) Where a+b+c+d+e=8 and a 2 +b 2 +c 2 +d 2 +e 2 =16 ∴ Eq. (i), reduces to ( 4 8−e ​ ) 2 ≤ 4 16−e 2 ​ ⇒ 64+e 2 −16e≤4(16−e 2 ) ⇒ 5e 2 −16e≤0 ⇒ e(5e−16)≤0 (Using number line rule) ⇒ 0≤e≤ 5 16 ​ Thus, range of eϵ[0, 5 16 ​ ]

cuzican:

does this help

OronSH:

this makes no sense, and it’s wrong

cuzican:

im just giving him a hand-

OronSH:

you’re writing nonsense and pretending it answers the question when it isn’t related at all

cuzican:

The point E can be obtained by folding the square along the bisector of angle BAC, so that AB is folded onto the diagonal with B landing on E. If F is the point where the bisector intersects B C, then B F folds to E F. Thus E F -1 AC and BF = EF. Since L.EFC = 45° = L.ECF, FE = EC. l.1(b). WC! = IBC! - IBFI = IAEI - ICEI = 21AEI - lAC! and IEC! = lAC! - IAEI = lAC! - IBC!. l.3(a). n Pn qn rn 1 1 1 1 2 3 2 3/2 = 1.5 3 7 5 7/5 = 1.4 4 17 12 17/12 = 1.416666 5 41 29 41/29 = 1.413793 1.3(c). Clearly, rn > 1 for each n, so that from (b), rn+1 - rn and rn - rn-I have opposite signs and 1 Irn+l - rnl < 41rn - rn-II· Since r2 > 1 = rl, it follows that rl < r3 < r2. Suppose as an induction hypothesis that rl < r3 < ... < r2m-1 < r2m < ... < r2. Then 0 < r2m - r2m+1 < r2m - r2m-1 and 0 < r2m+2 - r2m+1 < r2m - r2m+J. so that r2m-1 < r2m+1 < r2m+2 < r2m. Let k, 1 be any positive integers. Then, for m > k, I, r2k+1 < r2m-1 < r2m < r21. l.3(d). Leta be the least upper bound (i.e., the smallest number at least as great as all) of the numbers in the set {rl, r3, r5, ... , r2k+l, ... }. Then r2m-1 ::::: a ::::: r2m- the point is its useless

cuzican:

im using real math tho-

Extrinix:

Don’t you just love plagiarism

cuzican:

i cant-

OronSH:

don't you love it when someone copy pastes the solution to 1978 USAMO 1 instead of this problem

cuzican:

i really cant with yall- 🙃

cuzican:

I CANT-

OronSH:

please take the time to READ what you copy pasted. we're not dumb

cuzican:

Thick ya real smart saying I copy and pasted it?-

cuzican:

I cant-

OronSH:

well then actually solve the problem please

cuzican:

I Made A Joke.

cuzican:

I gave him wrong info for a reason

OronSH:

who is "him"

cuzican:

I said it already

cuzican:

The person?

cuzican:

Oh no-

OronSH:

bruh

cuzican:

IK YALL CANCELLED ME

cuzican:

-

cuzican:

did u use it..?

OronSH:

any actual help?

cuzican:

LOL.

cuzican:

Omg- I'm logging off for today.

Extrinix:

Okay, so in order to make \(\sf{~~~a^2+b^2+c^2+d^2+e^2+f^2}\) true, we need to find out what mixture of values will get you “\(\sf{\dfrac{36}{5}}\)” that also equal 6. So what values of a,b,c,d,e, and f, can not only get you “\(\sf{\dfrac{36}{5}}\),” but also get you 6?

OronSH:

i tried but i can't figure out how to make both equations true at the same time

Extrinix:

Try remembering negatives and decimals, I know it widens the range, but yk

OronSH:

it says positive though

OronSH:

so I plugged in a=b=c=d=1 and solved for e and f. I got \(1\pm\frac{\sqrt{15}}5\), which makes a^3+b^3+c^3+d^3+e^3+f^3 equal to \(\frac{48}5\) but how would I prove that it is the largest possible value?

Extrinix:

Go through the \(\sf{^3}\)s and solve it using the answers you got for the \(\sf{^2}\)s

OronSH:

how would i prove that it's the largest possible value though? it's possible there are different a, b, c, d, e, f that give a larger value

Extrinix:

Unless- was that the answer you got for the 3rds-?

Extrinix:

Well the possibilities of a, b, c, d, e, and f all have to equal 6 (as in the basic equation) But using the \(\sf{^2}\)s and solving what the values of a, b, c, d, e, and f are is what will give you the value for the \(\sf{^3}\)s

Extrinix:

If you get what I’m saying

OronSH:

there are infinitely many possibilities for a,b,c,d,e,f, how do i try all of them

Extrinix:

That’s why the squares are there

OronSH:

there's still infinitely many though

Extrinix:

For the first equation there is, but not if you solve for the variables

Extrinix:

On the second equation*

OronSH:

if i plug in any values of a,b,c,d then there are two variables left and two equations which gives a unique solution for e and f, so how do I try all possibilities of a,b,c,d? they are real numbers so there are infinitely many possibilities

Extrinix:

Look at it this way, what values can a b c d e and f equal that give you “36/5” when they are squared, but give you “6” when they are not squared?

OronSH:

infinitely many

Extrinix:

You have to get an exact value of 7.2 though.

OronSH:

there are 6 variables but only 2 equations, there's not enough information to solve for the variables

Extrinix:

You have to use the fact that a b c d e and f (their values themselves) all combine to equal 6 in every equation 36/5 is just the answer to what the variables all squared would equal

OronSH:

i found a few more: (0.8,0.8,0.8,0.8,0.8,2), (0.8,0.8,0.8,0.8,2,0.8), (0.8,0.8,0.8,2,0.8,0.8), (0.8,0.8,2,0.8,0.8,0.8),(0.8,2,0.8,0.8,0.8,0.8),(2,0.8,0.8,0.8,0.8,0.8), which all make the cubed expression 10.56, and (0,1.2,1.2,1.2,1.2,1.2), which makes the cubed expression 8.64

OronSH:

how do i prove that it is the maximum?

Extrinix:

Okay so I looked back to your answer that you got from solving them, and there’s an easier way to prove this, if I’m remembering this correctly The change between \(\sf{^1}\) and \(\sf{^2}\) is 1.2 Now, given that and the value you achieved, \(\sf{\dfrac{48}{5}}\), you can prove it This is simply by understanding the equation has a straight increase, so you can use the amount of increase per powers to figure out the next \(=\) It jumps from \(6\) to \(7.2\) (1.2 increase) and then to \(9.6\), which is a 2.4 increase Meaning your equation is increasing at this straight incline

OronSH:

but I found a larger value with different a, b, c, d, e, f, which is 10.56

minesweeper:

this is trivial

OronSH:

when did i ask

OronSH:

if it's trivial then what's your proof

OronSH:

ok I think I have a proof that 10.56 is optimal first b+c+d+e+f=6-a and b^2+c^2+d^2+e^2+f^2=36/5-a^2 so by cauchy schwarz (36-5a^2)>=(6-a)^2, so 6a^2-12a<=0, which means a<=2. Now, we have \((a-4/5)^2 (a-2)\leq0\) so \(a^3-3.6a^2+3.84a-1.28\leq0\), so \(a^3+b^3+c^3+d^3+e^3+f^3\\\leq3.6(a^2+b^2+c^2+d^2+e^2+f^2)-3.84(a+b+c+d+e+f)+7.68\\=25.92-23.04+7.68=10.56\) this means that the maximum value is \(\boxed{10.56}\) :)

cuzican:

nice -

cuzican:

u have time to do that but not google?

minesweeper:

alecks bad

OronSH:

@cuzican wrote:
u have time to do that but not google?
what do you mean google???

OronSH:

also oren sobad imagien only getting 35.7%